### 29 Mar 2026 v1

@@ Abstract
Real-world mathematical modeling is inherently an experiential and collaborative endeavor. Domain experts rarely solve complex problems from scratch; instead, they draw upon analogies from historical cases and subject their hypotheses to rigorous peer scrutiny. However, autonomous agents powered by Large Language Models predominantly rely on isolated reasoning paradigms, frequently generating plausible but fundamentally flawed models due to a lack of domain grounding and adversarial verification. To address these limitations, we propose Sci-Mind, a novel framework that mirrors the human scientific discovery process. Sci-Mind integrates Experiential Memory Recall to retrieve executable code snippets and modeling paradigm descriptors, grounding abstract reasoning in historical solutions. Subsequently, it employs an Adversarial Cognitive Dialectic where a Theorist optimizing mathematical coherence and a Pragmatist enforcing data feasibility debate through competing objectives to prune elegant but infeasible formulations. A Self-Validating Execution Strategy further ensures blueprint consistency through formal predicates before code generation, achieving fully autonomous execution. Extensive experiments on the MM-Bench and EngiBench benchmarks demonstrate that Sci-Mind significantly outperforms leading autonomous agents in both modeling rigorousness and code executability.
footnotetext:  Equal contribution; ✉ corresponding author.
@@ 1 Introduction
Mathematical modeling serves as the fundamental bridge translating complex physical phenomena into formal analytical systems [13]. Despite its wide applications in operations research [4], physical simulations [21], and biomedical network analysis [18, 19], autonomous tracking of open-ended scientific tasks relying solely on isolated reasoning often fails under complex real-world constraints [36]. In contrast, human scientific methodology has the advantage of experiential and collaborative frameworks, providing rigorous verification and historical grounding [35]. To integrate these advantages, augmenting Large Language Models (LLMs) with scientific reasoning paradigms is crucial for robust performance.
@@
However, current autonomous agents model the problem using zero-shot inference from the initial prompt, which often leads to severe theoretical drift or execution failures [20]. This limitation stems from two main factors. First, a single isolated reasoning process offers limited structural information about the target problem, failing to capture complex modeling paradigms across different scientific domains [31]. Second, the inherent ambiguity of real-world data often causes agents to overemphasize elegant but unvalidated theoretical features, resulting in practical infeasibility [45]. Moreover, directly translating abstract math into code without structural priors is highly brittle in practical scenarios [25], consequently hindering its widespread deployment. As illustrated in Fig. 1(a), these agents operate without experiential memory or adversarial verification, causing both execution failures and theoretical drift in the generated code.
@@
On the other hand, recent advances in LLM reasoning have focused on integrating self-reflection mechanisms [34]. In previous works, actor and critic models are first prompted independently, then fused through iterative feedback loops operating at the semantic level [37]. However, a significant portion of this self-reflection is redundant, forming an echo chamber that amplifies generative biases without exposing logical blind spots [38]. As shown in Fig. 1(b), when the generator and critic share the same objective, the resulting code may execute but carries reinforced biases that remain undetected. Furthermore, heterogeneous domain gaps between abstract theoretical logic and concrete programming syntax hinder the establishment of effective code execution [46]. This raises a basic question: How can we achieve deep theoretical rigor while escaping echo-chamber redundancies and bridging the gap to executable code?
@@
To address these challenges, we introduce cognitive bionics as a high-level representation to enhance mathematical modeling in LLMs. Human-like scientific discovery provides a more grounded understanding of problems than isolated reasoning, addressing structural limitations in abstract generation [26]. Experiential memory representation can offer clear semantic priors, including logic paradigms and executable demonstrations, enabling effective theory-implementation separation [24]. Recent advances in Retrieval-Augmented Generation (RAG) have demonstrated the potential of external knowledge for robust reasoning [27]. However, these methods face the challenge of semantic misalignment, as they predominantly fetch unstructured text that fails to map onto formal executable structures [3]. To bridge this gap, we extend current retrieval methods by introducing executable code-level insights and a non-cooperative game mechanism.
@@
With the above components, our Sci-Mind enables comprehensive scientific reasoning, leveraging both historical evidence and adversarial context to maintain theoretical rigor under complex scenarios. Extensive experiments on the MM-Bench [30] and EngiBench [48] datasets demonstrate the effectiveness of our method. In summary, our main contributions are as follows:
@@ 2 Related Work
2.1 Autonomous Agents for Scientific Reasoning
Autonomous agents powered by LLMs have emerged as a promising paradigm for decomposing and solving complex analytical tasks [40, 44]. Early frameworks leverage Zero-shot CoT prompting [41, 23] to elicit step-by-step reasoning, while DS-Agent [12] and MM-Agent [30] advance workflow standardization for data science and mathematical modeling respectively. In parallel, Coscientist [5] and ChemCrow [32] demonstrate the potential of LLM agents in autonomous experimental design. Recently, RAG [27] has been integrated into scientific agents to mitigate hallucinations [10], but existing approaches predominantly retrieve unstructured textual definitions that fail to bridge the gap between abstract formulations and executable code. To address this limitation, Sci-Mind introduces Experiential Memory Recall to retrieve executable code snippets alongside modeling paradigm descriptors, providing both procedural and conceptual priors for grounded reasoning.
@@ 2.2 Self-Reflection and Multi-Agent Debate
Self-reflection mechanisms enable LLMs to iteratively critique and refine their outputs. Reflexion [37] maintains a verbal memory of past failures to guide future attempts, while Self-Refine [33] prompts a single model to generate, critique, and revise in a closed loop. Multi-agent frameworks extend this idea by simulating collaborative environments: CAMEL [28] introduces inception prompting for role-playing communication, MetaGPT [15] assigns specialized software engineering roles to distinct agents, FetalAgents [16] demonstrates multi-agent collaboration for medical image analysis, and AutoGen [42] adapts the Actor-Critic paradigm for LLM-based hypothesis generation and evaluation. Multi-agent debate methods [9, 29, 6] further enhance robustness through interactive multi-round discussions. However, when all agents share the same implicit objective, the interaction frequently degenerates into an echo chamber that reinforces biases without exposing fundamental flaws [17]. This is particularly problematic in open-ended scientific modeling, where theoretical validity and data feasibility constitute structurally different evaluation dimensions that a single shared objective cannot simultaneously capture. In contrast, Sci-Mind introduces the Adversarial Cognitive Dialectic, in which a Theorist optimizing mathematical coherence and a Pragmatist enforcing data feasibility engage in structured debate with explicitly competing objectives.
@@
2.3 Code Generation and Autonomous Execution
Code generation agents translate natural language into executable scripts, with HumanEval [7] establishing the evaluation paradigm, LDB [47] advancing step-by-step runtime verification, Voyager [39] introducing self-improving code libraries, and Data Interpreter [14] decomposing data science problems via hierarchical graph modeling. Recent agentic systems [30, 48] further adopt sandboxed environments with automated error tracing. Despite these advances, direct logic-to-code translation remains brittle because modeling decisions and implementation decisions are entangled in a single generation pass. Our approach decouples these concerns through a Self-Validating Execution Strategy that interposes a structured JSON blueprint and an automated Verifier agent checking formal consistency predicates before code generation proceeds.
@@ 3 Methodology
In this paper, we propose Sci-Mind, a cognitively-inspired framework for autonomous mathematical modeling. As shown in Fig 2, Sci-Mind consists of three tightly integrated components: Experiential Memory Recall (EMR) for grounding abstract reasoning in executable structural priors, Adversarial Cognitive Dialectic (ACD) for jointly optimizing theoretical rigor and data feasibility through structured adversarial interaction, and Self-Validating Execution Strategy (SVE) for autonomous blueprint verification and robust code synthesis. We further introduce an Epistemic Self-Evolution mechanism to enable continual knowledge accumulation. Details are described as follows.
3.1 Overall Framework
At reasoning step $t$, the system receives a user query $\mathbf{Q}\in\mathbb{R}^{N\times d}$, where $N$ is the token length and $d$ is the hidden dimension. The overall pipeline can be formulated as a sequential composition of three mappings:
@@ 3.1 Overall Framework
where $\mathcal{R}$ denotes the EMR retrieval operator, $\mathcal{A}$ represents the ACD refinement process, and $\mathcal{S}$ is the SVE code synthesis module. The knowledge base $\mathcal{K}$ is maintained as a dynamic repository of structured triplets. The output $\mathbf{C}^{*}$ is an executable code artifact that is validated within a sandboxed environment $\mathcal{E}$.
@@ 3.2 Experiential Memory Recall (EMR)
A critical bottleneck in autonomous modeling is the abstraction-implementation gap: an agent may possess sufficient declarative knowledge to formulate a model but lack the procedural knowledge to implement it as executable code. Existing retrieval-augmented methods [22] predominantly fetch unstructured textual definitions, which provide semantic context but fail to supply the syntactic scaffolding necessary for code generation. To bridge this gap, EMR retrieves structured historical solutions as executable priors.
3.2.1 Multi-Tier Knowledge Representation.
@@
3.2.2 Domain-Aware Retrieval.
@@ 3.3 Adversarial Cognitive Dialectic (ACD)
Standard self-reflection methods prompt a single model to iteratively critique its own output. Recent studies [17] have shown that such homogeneous reflection frequently degenerates into an echo chamber that reinforces generative biases without exposing fundamental flaws. We argue that the root cause is objective homogeneity: both the generator and the critic optimize toward the same implicit objective. To address this, ACD decomposes the verification task into two explicitly competing objectives evaluated by functionally asymmetric agents, drawing on the principle that structured disagreement produces more robust outcomes than unchallenged consensus.
3.3.1 Dual-Objective Formulation.
@@
3.3.2 Structured Adversarial Interaction.
@@
$|\Gamma^{r}-\Gamma^{r-1}|<\epsilon\quad\text{and}\quad\min\{U_{T}(\mathbf{m}^{r}),\,U_{P}(\mathbf{m}^{r}\mid\mathcal{D}_{obs})\}>\gamma$ (12)
@@ 3.4 Self-Validating Execution Strategy (SVE)
3.4.1 Blueprint Construction.
The Architect module translates the equilibrium hypothesis $\mathbf{m}^{*}$ into a structured JSON blueprint $\mathcal{B}$, which serves as an intermediate representation between abstract modeling logic and concrete code. This design draws on the principle that separating specification from implementation reduces cascading errors in automated code synthesis [8]. Formally, $\mathcal{B}$ is defined as a tuple:
@@
3.4.2 Automated Consistency Verification.
@@
3.4.3 Self-Correcting Execution.
@@ 3.5 Epistemic Self-Evolution
To enable lifelong learning beyond a static knowledge base, we design a continuous self-evolution mechanism [39] that consolidates successful reasoning experiences into reusable knowledge.
3.5.1 Knowledge Distillation.
@@
3.5.2 Diversity-Aware Admission.
@@
where $\tau\in(0,1)$ is the novelty threshold. This mechanism ensures that the knowledge base grows in informational diversity rather than volume, maintaining retrieval efficiency while expanding the coverage of solvable problem types.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate on two complementary benchmarks. MM-Bench [30] contains 111 real-world mathematical modeling problems from MCM/ICM competitions spanning 2000 to 2025, covering 10 domains and 8 task types with end-to-end problem solving requirements. EngiBench Level 3 [48] provides 43 open-ended engineering modeling tasks from MCM/ICM, CUMCM, and APMCM spanning 2010 to 2024, covering Systems & Control, Physical & Structural, and Chemical & Biological subfields with official rubrics.
Baselines. We compare against seven baselines spanning direct prompting, specialized agents, and multi-agent frameworks. Direct Prompting applies Zero-shot CoT [41] to the backbone LLM. ReAct [44] interleaves reasoning with acting. DS-Agent [12] leverages case-based reasoning from Kaggle. CAMEL [28] enables role-playing multi-agent communication. AutoGen [42] supports flexible multi-agent conversation with code execution. MM-Agent [30] is the current state-of-the-art framework for mathematical modeling. Human Expert solutions are those awarded Honorable Mention or higher in actual MCM/ICM competitions.
Metrics. Following the MM-Bench evaluation protocol [30], we report four qualitative dimensions scored on a 1 to 10 scale: Analysis Evaluation (AE), Modeling Rigorousness (MR), Practicality (PS), and Result Analysis (RBA). The Avg. score is their arithmetic mean. We additionally report Code Executability (CE), defined as the percentage of generated solutions that execute without runtime errors. Each problem is evaluated 3 times and we report the mean with standard deviation. For EngiBench, we follow its official protocol [48].
4.2 Implementation Details
Backbones. To disentangle method-level contributions from backbone capabilities, we evaluate Sci-Mind and key baselines across four representative LLMs: GPT-4o [1], Claude-3.5-Sonnet [2], DeepSeek-R1 [11], and Qwen-2.5-72B [43]. These four backbones span proprietary and open-source models, general-purpose and reasoning-specialized architectures, and a range of parameter scales. All experiments use default temperature settings provided by the respective API providers. Multi-agent baselines are given equivalent computational budgets in terms of total LLM calls per problem to ensure fair comparison.
Hyperparameters. The EMR knowledge base $\mathcal{K}$ is constructed from 847 historical entries sourced from Kaggle competition solutions and open-access scientific code repositories, each processed into the tri-level format of problem embedding, verified code snippet, and paradigm descriptor. We use the top-$k=3$ retrieved entries per query. The dense encoder $\mathcal{T}$ in EMR uses a pre-trained Sentence-BERT model. For ACD, the balancing coefficient is set to $\lambda=0.5$, the quality threshold to $\gamma=0.75$, the convergence tolerance to $\epsilon=0.05$, and the maximum debate rounds to $R_{max}=6$. For SVE, the blueprint revision budget is $T_{max}=3$ and the code retry budget is $J_{max}=5$. For self-evolution, the novelty threshold is $\tau=0.85$.
4.3 Main Results
@@
4.4 Cross-Domain Generalization on EngiBench

### 02 Apr 2026 v2

@@ Abstract
Real-world mathematical modeling is inherently an experiential and collaborative endeavor. Domain experts rarely solve complex problems from scratch; instead, they draw upon analogies from historical cases and subject their hypotheses to rigorous peer scrutiny. However, autonomous agents powered by Large Language Models predominantly rely on isolated reasoning paradigms, frequently generating plausible but fundamentally flawed models due to a lack of domain grounding and adversarial verification. To address these limitations, we propose Sci-Mind, a novel framework that mirrors the human scientific discovery process. Sci-Mind integrates Experiential Memory Recall to retrieve executable code snippets and modeling paradigm descriptors, grounding abstract reasoning in historical solutions. Subsequently, it employs an Adversarial Cognitive Dialectic where a Theorist optimizing mathematical coherence and a Pragmatist enforcing data feasibility debate through competing objectives to prune elegant but infeasible formulations. A Self-Validating Execution Strategy further ensures blueprint consistency through formal predicates before code generation, achieving fully autonomous execution. Extensive experiments on the MM-Bench and EngiBench demonstrate that Sci-Mind significantly outperforms leading autonomous agents in both modeling rigorousness and code executability.
@@ 1 Introduction
Mathematical modeling serves as the fundamental bridge translating complex physical phenomena into formal analytical systems Hao et al. (2024). Despite its wide applications in operations research Bengio et al. (2021), physical simulations Karniadakis et al. (2021), and biomedical network analysis Jia et al. (2025a, b), autonomous tracking of open-ended scientific tasks relying solely on isolated reasoning often fails under complex real-world constraints Romera-Paredes et al. (2024). In contrast, human scientific methodology has the advantage of experiential and collaborative frameworks, providing rigorous verification and historical grounding Popper (2005). To integrate these advantages, augmenting Large Language Models (LLMs) with scientific reasoning paradigms is crucial for robust performance.
@@
However, current autonomous agents model the problem using zero-shot inference from the initial prompt, which often leads to severe theoretical drift or execution failures Jimenez et al. (2024). This limitation stems from two main factors. First, a single isolated reasoning process offers limited structural information about the target problem, failing to capture complex modeling paradigms across different scientific domains Lu et al. (2024). Second, the inherent ambiguity of real-world data often causes agents to overemphasize elegant but unvalidated theoretical features, resulting in practical infeasibility Yin et al. (2024). Moreover, directly translating abstract math into code without structural priors is highly brittle in practical scenarios Lai et al. (2023), consequently hindering its widespread deployment. As illustrated in Fig. 1(a), these agents operate without experiential memory or adversarial verification, causing both execution failures and theoretical drift in the generated code.
@@
On the other hand, recent advances in LLM reasoning have focused on integrating self-reflection mechanisms Pan et al. (2024). In previous works, actor and critic models are first prompted independently, then fused through iterative feedback loops operating at the semantic level Shinn et al. (2023). However, a significant portion of this self-reflection is redundant, forming an echo chamber that amplifies generative biases without exposing logical blind spots Valmeekam et al. (2023). As shown in Fig. 1(b), when the generator and critic share the same objective, the resulting code may execute but carries reinforced biases that remain undetected. Furthermore, heterogeneous domain gaps between abstract theoretical logic and concrete programming syntax hinder the establishment of effective code execution Zan et al. (2023). This raises a basic question: How can we achieve deep theoretical rigor while escaping echo-chamber redundancies and bridging the gap to executable code?
@@
To address these challenges, we introduce cognitive bionics as a high-level representation to enhance mathematical modeling in LLMs. Human-like scientific discovery provides a more grounded understanding of problems than isolated reasoning, addressing structural limitations in abstract generation Langley (1987). Experiential memory representation can offer clear semantic priors, including logic paradigms and executable demonstrations, enabling effective theory-implementation separation Kolodner (1992). Recent advances in Retrieval-Augmented Generation (RAG) have demonstrated the potential of external knowledge for robust reasoning Lewis et al. (2020). However, these methods face the challenge of semantic misalignment, as they predominantly fetch unstructured text that fails to map onto formal executable structures Barnett et al. (2024). To bridge this gap, we extend current retrieval methods by introducing executable code-level insights and a non-cooperative game mechanism.
@@
With the above components, our Sci-Mind enables comprehensive scientific reasoning, leveraging both historical evidence and adversarial context to maintain theoretical rigor under complex scenarios. Extensive experiments on the MM-Bench Liu et al. (2025) and EngiBench Zhou et al. (2025) datasets demonstrate the effectiveness of our method. In summary, our main contributions are as follows:
@@ 2 Related Work
2.1 Scientific Reasoning Agents
LLM-powered autonomous agents have become a prominent paradigm for complex analytical tasks Wang et al. (2024b); Yao et al. (2023), progressing from Zero-shot CoT prompting Wei et al. (2022); Kojima et al. (2022) to specialized workflows such as DS-Agent Guo et al. (2024) for data science and MM-Agent Liu et al. (2025) for mathematical modeling, as well as domain-specific scientific agents like Coscientist Boiko et al. (2023) and ChemCrow M. Bran et al. (2024). While integrating RAG Lewis et al. (2020) helps mitigate hallucinations Gao et al. (2023), existing retrieval methods predominantly fetch unstructured text that fails to bridge abstract formulations and executable code. Sci-Mind addresses this gap through Experiential Memory Recall, which retrieves executable code snippets alongside modeling paradigm descriptors, providing both procedural and conceptual priors.
@@ 2.2 Self-Reflection and Multi-Agent Debate
Self-reflection methods such as Reflexion Shinn et al. (2023) and Self-Refine Madaan et al. (2023) enable iterative critique within a single model, while multi-agent frameworks, including CAMEL Li et al. (2023), MetaGPT Hong et al. (2024), FetalAgents Hu et al. (2026), and AutoGen Wu et al. (2024), extend this idea through collaborative role-based interaction. Multi-agent debate Du et al. (2024); Liang et al. (2024); Chan et al. (2024) further improves robustness via multi-round discussion. However, when all agents share the same implicit objective, the interaction tends to degenerate into an echo chamber that reinforces biases rather than exposing fundamental flaws Huang et al. (2024). This is especially problematic in scientific modeling, where theoretical validity and data feasibility are structurally distinct evaluation dimensions. Sci-Mind addresses this through the Adversarial Cognitive Dialectic, in which a Theorist and a Pragmatist debate with explicitly competing objectives.
2.3 Code Generation and Execution
Code generation research spans evaluation benchmarks Chen et al. (2021), runtime verification Zhong et al. (2024), self-improving code libraries Wang et al. (2024a), hierarchical task decomposition Hong et al. (2025), and sandboxed agentic execution Liu et al. (2025); Zhou et al. (2025). Despite these advances, direct logic-to-code translation remains brittle when modeling and implementation decisions are entangled in a single generation pass. Sci-Mind decouples these concerns through a Self-Validating Execution Strategy that interposes a structured JSON blueprint and an automated Verifier with formal consistency predicates before code generation.
@@ 3 Methodology
In this paper, we propose Sci-Mind, a cognitively-inspired framework for autonomous mathematical modeling, as shown in Fig 2.
Formally, given a user query $\mathbf{Q}\in\mathbb{R}^{N\times d}$ where $N$ is the token length and $d$ is the hidden dimension, the overall pipeline is formulated as a sequential composition of three mappings:
@@
where $\mathcal{R}$ denotes the EMR retrieval operator, $\mathcal{A}$ represents the ACD refinement process, and $\mathcal{S}$ is the SVE code synthesis module. The knowledge base $\mathcal{K}$ is maintained as a dynamic repository of structured triplets, and the output $\mathbf{C}^{*}$ is an executable code artifact validated within a sandboxed environment $\mathcal{E}$. Details of each component are described as follows.
@@ 3.1 Experiential Memory Recall (EMR)
A critical bottleneck in autonomous modeling is the abstraction-implementation gap: an agent may possess sufficient declarative knowledge to formulate a model but lack the procedural knowledge to implement it as executable code. Existing retrieval-augmented methods Karpukhin et al. (2020) predominantly fetch unstructured textual definitions, which provide semantic context but fail to supply the syntactic scaffolding necessary for code generation. To bridge this gap, EMR retrieves structured historical solutions as executable priors.
3.1.1 Multi-Tier Knowledge Representation
@@
3.1.2 Domain-Aware Retrieval
@@ 3.2 Adversarial Cognitive Dialectic (ACD)
Standard self-reflection methods prompt a single model to iteratively critique its own output. Recent studies Huang et al. (2024) have shown that such homogeneous reflection frequently degenerates into an echo chamber that reinforces generative biases without exposing fundamental flaws. We argue that the root cause is objective homogeneity: both the generator and the critic optimize toward the same implicit objective. To address this, ACD decomposes the verification task into two explicitly competing objectives evaluated by functionally asymmetric agents, drawing on the principle that structured disagreement produces more robust outcomes than unchallenged consensus.
3.2.1 Dual-Objective Formulation
@@
3.2.2 Structured Adversarial Interaction
@@
$|\Gamma^{r}-\Gamma^{r-1}|<\epsilon\quad\text{and}$ (12)
$\min\{U_{T}(\mathbf{m}^{r}),\,U_{P}(\mathbf{m}^{r}\mid\mathcal{D}_{obs})\}>\gamma$
@@ 3.3 Self-Validating Execution Strategy (SVE)
3.3.1 Blueprint Construction
The Architect module translates the equilibrium hypothesis $\mathbf{m}^{*}$ into a structured JSON blueprint $\mathcal{B}$, which serves as an intermediate representation between abstract modeling logic and concrete code. This design draws on the principle that separating specification from implementation reduces cascading errors in automated code synthesis Chen et al. (2024). Formally, $\mathcal{B}$ is defined as a tuple:
@@ 3.3.1 Blueprint Construction
3.3.2 Automated Consistency Verification
@@
3.3.3 Self-Correcting Execution
@@ 3.4 Epistemic Self-Evolution
To enable lifelong learning beyond a static knowledge base, we design a continuous self-evolution mechanism Wang et al. (2024a) that consolidates successful reasoning experiences into reusable knowledge.
3.4.1 Knowledge Distillation
@@
3.4.2 Diversity-Aware Admission
@@
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate on two complementary benchmarks. MM-Bench Liu et al. (2025) contains 111 real-world mathematical modeling problems from MCM/ICM competitions spanning 2000 to 2025, covering 10 domains and 8 task types with end-to-end problem solving requirements. EngiBench Level 3 Zhou et al. (2025) provides 43 open-ended engineering modeling tasks from MCM/ICM, CUMCM, and APMCM spanning 2010 to 2024, covering Systems & Control, Physical & Structural, and Chemical & Biological subfields with official rubrics.
Baselines. We compare against seven baselines spanning direct prompting, specialized agents, and multi-agent frameworks. Direct Prompting applies Zero-shot CoT Wei et al. (2022) to the backbone LLM. ReAct Yao et al. (2023) interleaves reasoning with acting. DS-Agent Guo et al. (2024) leverages case-based reasoning from Kaggle. CAMEL Li et al. (2023) enables role-playing multi-agent communication. AutoGen Wu et al. (2024) supports flexible multi-agent conversation with code execution. MM-Agent Liu et al. (2025) is the current state-of-the-art framework for mathematical modeling. Human Expert solutions are those awarded Honorable Mention or higher in actual MCM/ICM competitions.
Backbones. We evaluate Sci-Mind and key baselines across four representative LLMs: GPT-4o Achiam et al. (2023), Claude-3.5-Sonnet Anthropic (2024), DeepSeek-R1 Guo et al. (2025), and Qwen-2.5-72B Yang et al. (2024).
Metrics. Following the MM-Bench evaluation protocol Liu et al. (2025), we report four qualitative dimensions scored on a 1 to 10 scale: Analysis Evaluation (AE), Modeling Rigorousness (MR), Practicality (PS), and Result Analysis (RBA). The Avg. score is their arithmetic mean. We additionally report Code Executability (CE), defined as the percentage of generated solutions that execute without runtime errors. Each problem is evaluated 3 times and we report the mean with standard deviation. For EngiBench, we follow its official rubric-based protocol Zhou et al. (2025).